OpenStudy (anonymous):

If y varies inversely with x and y = 3 when x = 4, what is the value of y when x = 2?

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \bf y=\frac{k}{x} }\) is the inverse variation form.

OpenStudy (anonymous):

A. \[y=\frac{ 1 }{ 6 }\] B. y=2 C. y=6 D. y=12

OpenStudy (anonymous):

@SolomonZelman Thanks you

OpenStudy (solomonzelman):

these are the options, as I can see.... \(\Large\color{red}{ \bf y=\frac{k}{x} }\) plugging in 4 for x and 3 for y, \(\Large\color{red}{ \bf 3=\frac{k}{4} }\) multiply both sides by 4, and k=12. NOW, .... when x=2, y=? \(\Large\color{red}{ \bf ?=\frac{12}{2} }\)